# A Comprehensive Survey of Mixture-of-Experts: Algorithms, Theory, and Applications

### v1

@@ Abstract
Artificial intelligence (AI) has achieved astonishing successes in many domains, especially with the recent breakthroughs in the development of foundational large models. These large models, leveraging their extensive training data, provide versatile solutions for a wide range of downstream tasks. However, as modern datasets become increasingly diverse and complex, the development of large AI models faces two major challenges: (1) the enormous consumption of computational resources and deployment difficulties, and (2) the difficulty in fitting heterogeneous and complex data, which limits the usability of the models. Mixture of Experts (MoE) models has recently attracted much attention in addressing these challenges, by dynamically selecting and activating the most relevant sub-models to process input data. It has been shown that MoEs can significantly improve model performance and efficiency with fewer resources, particularly excelling in handling large-scale, multimodal data. Given the tremendous potential MoE has demonstrated across various domains, it is urgent to provide a comprehensive summary of recent advancements of MoEs in many important fields. Existing surveys on MoE have their limitations, e.g., being outdated or lacking discussion on certain key areas, and we aim to address these gaps. In this paper, we first introduce the basic design of MoE, including gating functions, expert networks, routing mechanisms, training strategies, and system design. We then explore the algorithm design of MoE in important machine learning paradigms such as continual learning, meta-learning, multi-task learning, and reinforcement learning. Additionally, we summarize theoretical studies aimed at understanding MoE and review its applications in computer vision and natural language processing. Finally, we discuss promising future research directions.
@@
Mixture-of-Experts, Continual Learning, Meta-Learning, Multi-Task Learning, Reinforcement Learning, Computer Vision, Natural Language Processing
@@ I Introduction
for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=line-color, rounded corners,align=left, minimum width=4.0em, s sep=4pt, inner xsep=4pt, inner ysep=4pt, align=left, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=5.5em,font=,, where level=2text width=5.5em,font=, where level=3text width=5.5em,font=, where level=4text width=5.5em,font=, [Mixture of Experts (MoE), ver [Basics of MoE [Gating Function [Switch transformers [15], V-MoE [17], RMoE [18], M3vit [19], GLaM [20], Shazeer et al. [16],
@@
PMoE [34], Wang et al. [68], Chen et al. [69], Wang et al. [70] , Park et al. [64], Aljundi et al. [71] ,
Wang et al. [72] , Doan et al. [73] ,leaf,text width=30em] ] [Meta
@@
TERL [126] , ,leaf,text width=30em] ] ] [Theory [Nguyen et al. [25], Chen et al. [29], Li et al. [30], Li et al. [30], Chowdhury et al. [127],Jiang et al. [128], Jiang et al. [129],
@@
This paper will be organized as follows (as illustrated in Figure 1). In Section II, we will provide a comprehensive introduction to the basic designs of MoE, ranging from basic architecture design to training strategies to system optimization strategies. In Section III, we will introduce the recent MoE-based algorithm design in four important machine learning paradigms, including continual learning, meta-learning, multi-task learning, and reinforcement learning, which seeks to provide readers a general idea about how MoE can be leveraged to improve the algorithm design in these domains. In Section IV, we will summarize the efforts towards building the theoretical understanding of MoE, which covers a variety of aspects in MoE, such as different gating functions, different expert models, and different learning scenarios. In Section V, we will present the recent applications of MoE in two major domains, i.e., computer vision and natural language processing, where we look into multiple important subproblems in each domain. Future research directions will be discussed in Section VI, followed by the conclusion in Section VII.
@@ II-B Expert Networks
Compared to the standard multi-head attention mechanism [197] , the key advantage of MoA lies in its use of a gating network to dynamically select the most relevant attention heads for each input token and compute a weighted sum of their outputs. This approach eliminates the requirement for all attention heads to process every input, enabling MoA to achieve superior model performance with reduced computational overhead. Consequently, MoA exhibits enhanced scalability, and particularly this model can attain higher BLEU scores than Transformer with 213 million parameters while requiring significantly less computation.
Inspired by MoA, numerous studies have been proposed to improve upon this design. SwitchHead [31] attempts to independently introduce the MoE mechanism for each head’s key, query, value, and output projection, and employs a non-competitive selection activation function. This will therefore reduce the number of attention matrices that need to be computed, significantly lowering computational and memory requirements. MoH [33] enhances the standard MoE method by sharing heads and a two-stage routing strategy, and can further fine-tune pre-trained multi-head attention models into MoH models, enhancing their applicability. In addition, this work integrates the proposed MoH framework with various model frameworks including the encoder-decoder architecture, such as ViT [1], DiT [201], and decoder-only LLMs [202],
which are applied in various tasks and demonstrate great potentials of MoH.
@@ II-C Routing Strategy
2) Modality-Level Routing Strategy routes tokens based on their modality, emulating the practice of employing specialized encoders for distinct modalities to minimize interference between them. The Uni-MoE model [43] is designed to effectively process and understand multi-modal data through a structured approach. First, it introduces connectors trained using cross-modal data to achieve data alignment, mapping diverse modalities into a unified language representation space. Simultaneously, it trains modality-specific expert systems on cross-modal instruction datasets, so as to enhance each expert’s specialized capabilities and activate their task-specific preferences. Subsequently, the model utilizes LoRA [203]
for fine-tuning, optimizing performance with mixed multi-modal instruction data while maintaining computational efficiency. Following instruction tuning, Uni-MoE is evaluated on a comprehensive and diverse multi-modal dataset. The results indicate that the model significantly reduces performance discrepancies when processing mixed data, improves collaboration efficiency among multiple experts, and enhances generalization capabilities, highlighting the potential of the MoE architecture in developing a unified multi-modal large language model.
@@ II-D Training Strategies
Due to the sparse activations inherent in MoE-based architectures, directly applying the functions and procedures used for training dense models to MoE-based sparse models is not straightforward. To improve training effectiveness, achieve faster convergence, and enhance performance on target tasks, numerous studies like [15, 17, 16, 18, 46]
have explored the methods for training these sparse models. In the following section, we summarize and discuss some representative techniques.
@@
As mentioned before, BPR [17] achieves sparsity reuse by prioritizing important samples for processing. The advantages of BPR lie in its efficient resource utilization and reduction of unnecessary computations.
RMoE  [18] designs a pipeline particularly for downstream tasks such as segmentation and detection. More specifically, RMoE decomposes the weights of MoE experts into core and residual weights, where the core weights inherit knowledge from pre-trained non-MoE models and only the residual weights will be finetuned. This will significantly shorten the training time and reduce the training costs. During inference, both sets of weights will be combined to ensure the good performance of the MoE model.
@@ II-E System Design
The MoE architecture is gaining popularity in the edge deployment and capacity scaling of large models [15, 51, 52],
offering significant flexibility and scalability. However, its nature of inherent non-uniform workload distribution and dynamic expert selection imposes strict requirements on the system’s resource allocation, communication efficiency, and overall performance optimization. This will introduce new system challenges such as increased synchronization overhead, hampering the efficiency and stability of model execution. In this section, we outline some common methods currently available to enhance the system designs for MoE from three dimensions: computation, communication, and memory.
@@ III Algorithms
Due to the effectiveness and the strength in handling diverse knowledge in data, there have been rapidly growing interests in applying the MoE architecture into various machine learning paradigms, leading to improved algorithm designs. To showcase this and the great potentials of MoE in solving more complex learning scenarios, we will delve into the MoE-based algorithm designs for four widely investigated domains, i.e., continual learning, meta-learning, multi-task learning, and reinforcement learning.
@@ III-C Multi-task Learning
Multi-task reinforcement learning is an important branch of multi-task learning [225]. It aims to promote the learning of diverse skills by agents to enhance their ability to handle complex tasks. However, existing methods lack guarantees on the diversity of learned representations, which may lead to representation collapse into similar representations, thus limiting their generalization ability [96, 226, 96, 227, 228, 229]. To address this issue, [80] introduces a method called Mixture of Orthogonal Experts (MOOR), which utilizes multiple expert models responsible for generating different representations. In this case, each expert is responsible for different aspects of the task, encouraging the model to learn more diverse skills. To ensure efficient differentiations among experts, MOOR introduces the Gram-Schmidt orthogonalization method into MoE to force each expert to generate mutually orthogonal features, further limiting the possibility of experts generating redundant representations. [96] investigates contextual multi-task RL and proposed a novel approach to encode an input observation into multiple representations using a mixture of encoders. These representations correspond to different skills or objects, and can be selected for any given task based on the context information. This approach gives the learning agent a more fine-grained control on the information shared across tasks and alleviates negative interference. Instead of naively sharing parameters across all tasks in traditional approaches, [85] proposes an attention-based MoE approach for multi-task reinforcement learning, which seeks to learn a compositional policy for each task. In particular, each expert network learns task-specific skills and specializes in different parts of the multi-task representation space. An attention module was fu introduced to integrate the expert outputs, we propose an attention module to generate connections between tasks and experts to achieve the best performance automatically.
@@ IV Theory
Although the MoE model has a long history and achieved achieved great empirical success in practice during the past few years [248, 16], the theory behind it is largely under-explored, especially in deep neural networks. Recently, some studies have initiated the attempt to build theoretical understandings of MoE. In this section, we will delve into these studies, aiming to provide the readers a basic idea of the theory development in this area.
@@
With the increasingly broad applications of MoE in many problems, the theoretical studies of MoE start to emerge beyond the general machine learning setup. [139] explores the usage of MoE models in handling multilevel data while building the theoretical understanding therein. Notably, multilevel data is very common in practical applications [249, 250], such as hierarchical data with a nested structure where different levels are correlated. Ignoring the dependencies within the data can lead to spurious, misleading, or biased clustering and prediction outcomes [251]. To better capture the dependencies, [139] proposes a Mixed MoE (MMoE) architecture to handle multi-level data. This approach not only introduces random effects into the gating function, making the computed gating values depend not only on input variables but also on random effects. This helps the model better capture random effects in multi-level data when dynamically adjusting expert weights. At the same time, to simplify the model, the expert functions are separated from the random effects, and the probability distribution of the expert model no longer depends on input variables but only on output variables. Under certain regularity conditions, this paper proves that the MMoE is dense in the sense of weak convergence for any continuous mixed-effects model. In the special case of hierarchical multi-level data, this paper further proves that MMoE can approximate the complex dependency structures of random effects between different factor levels in multi-level data.
@@ V Applications
In this section, we investigate the applications of MoE in two broad application domains, i.e., computer vision (CV) and natural language processing (NLP), among which various specific applications will be explored.

### v2

@@ Abstract
Artificial intelligence (AI) has achieved astonishing successes in many domains, especially with the recent breakthroughs in the development of foundational large models. These large models, leveraging their extensive training data, provide versatile solutions for a wide range of downstream tasks. However, as modern datasets become increasingly diverse and complex, the development of large AI models faces two major challenges: (1) the enormous consumption of computational resources and deployment difficulties, and (2) the difficulty in fitting heterogeneous and complex data, which limits the usability of the models. Mixture of Experts (MoE) models has recently attracted much attention in addressing these challenges, by dynamically selecting and activating the most relevant sub-models to process input data. It has been shown that MoEs can significantly improve model performance and efficiency with fewer resources, particularly excelling in handling large-scale, multimodal data. Given the tremendous potential MoE has demonstrated across various domains, it is urgent to provide a comprehensive summary of recent advancements of MoEs in many important fields. Existing surveys on MoE have their limitations, e.g., being outdated or lacking discussion on certain key areas, and we aim to address these gaps. In this paper, we first introduce the basic design of MoE, including gating functions, expert networks, routing mechanisms, training strategies, and system design. We then explore the algorithm design of MoE in important machine learning paradigms such as continual learning, meta-learning, multi-task learning, reinforcement learning, and federated learning. Additionally, we summarize theoretical studies aimed at understanding MoE and review its applications in computer vision and natural language processing. Finally, we discuss promising future research directions.
@@
Mixture-of-Experts, Continual Learning, Meta-Learning, Multi-Task Learning, Reinforcement Learning, Federated Learning, Computer Vision, Natural Language Processing
@@ I Introduction
for tree= grow=east, reversed=true, anchor=base west, transform shape, scale=1.1, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=line-color, rounded corners,align=left, minimum width=4.0em, s sep=4pt, inner xsep=4pt, inner ysep=4pt, align=left, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=5.5em,font=,, where level=2text width=5.5em,font=, where level=3text width=5.5em,font=, where level=4text width=5.5em,font=, [Mixture of Experts (MoE), ver [Basics of MoE [Gating Function [Switch transformers [15], V-MoE [17], RMoE [18], M3vit [19], GLaM [20], Shazeer et al. [16],
@@
PMoE [34], Wang et al. [68], Chen et al. [69], Wang et al. [70] , Aljundi et al. [71] , Wang et al. [72],
Doan et al. [73] ,leaf,text width=30em] ] [Meta
@@
TERL [126] , ,leaf,text width=30em] ] [Federated
Learning [Peterson et al. [127], Zec et al. [128], Pye et al. [129], Reisser et al. [130], Guo et al. [131],
Isaksson et al. [132], Ghosh et al. [133], Tran et al. [134], Dun et al. [135], Heinbaught et al. [136],
Su et al. [137], Zeng et al. [138],leaf,text width=30em] ] ] [Theory [Nguyen et al. [25], Chen et al. [29], Li et al. [30], Li et al. [139], Chowdhury et al. [140],Jiang et al. [141], Jiang et al. [142],
@@
This paper will be organized as follows (as illustrated in Figure 1). In Section II, we will provide a comprehensive introduction to the basic designs of MoE, ranging from basic architecture design to training strategies to system optimization strategies. In Section III, we will introduce the recent MoE-based algorithm design in four important machine learning paradigms, including continual learning, meta-learning, multi-task learning, reinforcement learning, and federated learning, which seeks to provide readers a general idea about how MoE can be leveraged to improve the algorithm design in these domains. In Section IV, we will summarize the efforts towards building the theoretical understanding of MoE, which covers a variety of aspects in MoE, such as different gating functions, different expert models, and different learning scenarios. In Section V, we will present the recent applications of MoE in two major domains, i.e., computer vision and natural language processing, where we look into multiple important subproblems in each domain. Future research directions will be discussed in Section VI, followed by the conclusion in Section VII.
@@ II-B Expert Networks
Compared to the standard multi-head attention mechanism [210], the key advantage of MoA lies in its use of a gating network to dynamically select the most relevant attention heads for each input token and compute a weighted sum of their outputs. This approach eliminates the requirement for all attention heads to process every input, enabling MoA to achieve superior model performance with reduced computational overhead. Consequently, MoA exhibits enhanced scalability, and particularly this model can attain higher BLEU scores [302] than Transformer with 213 million parameters while requiring significantly less computation.
Inspired by MoA, numerous studies have been proposed to improve upon this design. SwitchHead [31] attempts to independently introduce the MoE mechanism for each head’s key, query, value, and output projection, and employs a non-competitive selection activation function. This will therefore reduce the number of attention matrices that need to be computed, significantly lowering computational and memory requirements. MoH [33] enhances the standard MoE method by sharing heads and a two-stage routing strategy, and can further fine-tune pre-trained multi-head attention models into MoH models, enhancing their applicability. In addition, this work integrates the proposed MoH framework with various model frameworks including the encoder-decoder architecture, such as ViT [1], DiT [214], and decoder-only LLMs [215], which are applied in various tasks and demonstrate great potentials of MoH.
@@ II-C Routing Strategy
2) Modality-Level Routing Strategy routes tokens based on their modality, emulating the practice of employing specialized encoders for distinct modalities to minimize interference between them. The Uni-MoE model [43] is designed to effectively process and understand multi-modal data through a structured approach. First, it introduces connectors trained using cross-modal data to achieve data alignment, mapping diverse modalities into a unified language representation space. Simultaneously, it trains modality-specific expert systems on cross-modal instruction datasets, so as to enhance each expert’s specialized capabilities and activate their task-specific preferences. Subsequently, the model utilizes LoRA [216] for fine-tuning, optimizing performance with mixed multi-modal instruction data while maintaining computational efficiency. Following instruction tuning, Uni-MoE is evaluated on a comprehensive and diverse multi-modal dataset. The results indicate that the model significantly reduces performance discrepancies when processing mixed data, improves collaboration efficiency among multiple experts, and enhances generalization capabilities, highlighting the potential of the MoE architecture in developing a unified multi-modal large language model.
@@ II-D Training Strategies
Due to the sparse activations inherent in MoE-based architectures, directly applying the functions and procedures used for training dense models to MoE-based sparse models is not straightforward. To improve training effectiveness, achieve faster convergence, and enhance performance on target tasks, numerous studies like [15, 17, 16, 18, 46] have explored the methods for training these sparse models. In the following section, we summarize and discuss some representative techniques.
@@
As mentioned before, BPR [17] achieves sparsity reuse by prioritizing important samples for processing. The advantages of BPR lie in its efficient resource utilization and reduction of unnecessary computations. RMoE [18] designs a pipeline particularly for downstream tasks such as segmentation and detection. More specifically, RMoE decomposes the weights of MoE experts into core and residual weights, where the core weights inherit knowledge from pre-trained non-MoE models and only the residual weights will be finetuned. This will significantly shorten the training time and reduce the training costs. During inference, both sets of weights will be combined to ensure the good performance of the MoE model.
@@ II-E System Design
The MoE architecture is gaining popularity in the edge deployment and capacity scaling of large models [15, 51, 52], offering significant flexibility and scalability. However, its nature of inherent non-uniform workload distribution and dynamic expert selection imposes strict requirements on the system’s resource allocation, communication efficiency, and overall performance optimization. This will introduce new system challenges such as increased synchronization overhead, hampering the efficiency and stability of model execution. In this section, we outline some common methods currently available to enhance the system designs for MoE from three dimensions: computation, communication, and memory.
@@ III Algorithms
Due to the effectiveness and the strength in handling diverse knowledge in data, there have been rapidly growing interests in applying the MoE architecture into various machine learning paradigms, leading to improved algorithm designs. To showcase this and the great potentials of MoE in solving more complex learning scenarios, we will delve into the MoE-based algorithm designs for four widely investigated domains, i.e., continual learning, meta-learning, multi-task learning, reinforcement learning, and federated learning.
@@ III-C Multi-task Learning
Multi-task reinforcement learning is an important branch of multi-task learning [238]. It aims to promote the learning of diverse skills by agents to enhance their ability to handle complex tasks. However, existing methods lack guarantees on the diversity of learned representations, which may lead to representation collapse into similar representations, thus limiting their generalization ability [96, 239, 240, 241, 242]. To address this issue, [80] introduces a method called Mixture of Orthogonal Experts (MOOR), which utilizes multiple expert models responsible for generating different representations. In this case, each expert is responsible for different aspects of the task, encouraging the model to learn more diverse skills. To ensure efficient differentiations among experts, MOOR introduces the Gram-Schmidt orthogonalization method into MoE to force each expert to generate mutually orthogonal features, further limiting the possibility of experts generating redundant representations. [96] investigates contextual multi-task RL and proposed a novel approach to encode an input observation into multiple representations using a mixture of encoders. These representations correspond to different skills or objects, and can be selected for any given task based on the context information. This approach gives the learning agent a more fine-grained control on the information shared across tasks and alleviates negative interference. Instead of naively sharing parameters across all tasks in traditional approaches, [85] proposes an attention-based MoE approach for multi-task reinforcement learning, which seeks to learn a compositional policy for each task. In particular, each expert network learns task-specific skills and specializes in different parts of the multi-task representation space. An attention module was fu introduced to integrate the expert outputs, we propose an attention module to generate connections between tasks and experts to achieve the best performance automatically.
@@ III-D Reinforcement learning
III-E Federated learning
Federated Learning [261] is proposed to address the challenge of how different data owners can effectively exchange data without compromising user privacy and ensuring data security. By first training multiple local models on different servers and then aggregating these local model parameters using various rules to update the global model, the federated learning paradigm not only fully leverages distributed yet information-rich data but also efficiently utilizes the computational resources of multiple clients, inspiring a series of subsequent work, particularly in the field of recommendation systems [262, 263, 264].
When updating the global model using multiple local model parameters, one of the most common rules is Federated Averaging (FedAVG) [265]. This rule assigns weights to the corresponding local models based on the size of each client’s local dataset during parameter updates, and then updates the global model using the weighted combinations of local parameters. However, this rule relies on a manually defined simple heuristic, which not only requires each model to be initialized with the same random seed for convergence but is also unsuitable for scenarios where the data distribution among clients is highly heterogeneous [265]. Considering the hybrid nature of MoE, [127] is the first to introduce the MoE architecture into the federated learning setting. Specifically, this work proposes to train a general model using aggregated data and domain-specific models for each client using private data, and then use a gating mechanism to determine the contribution of each model’s output to the learning task. Compared to FedAVG, this approach better generates outputs based on task characteristics. [128] adopts a similar approach but focuses more on scenarios with stronger data heterogeneity across clients. Additionally, an opt-out mechanism is designed to ensure that clients with sensitive data could choose not to participate in the data integration process, thereby safeguarding data security. Some other work has also followed this combined approach [129].
An important goal of federated learning is to effectively integrate data from multiple clients. However, it is evident that real-world data distributions are highly complex and variable. Therefore, a major challenge in federated learning is data heterogeneity, which can manifest in two ways: one where clients share the same conditional distribution but have different marginal distributions, and the other where the opposite is true. Both scenarios can harm the model’s learning performance. Recently some studies have made substantial efforts to address this challenge. For example, [130] introduces two improvements: First, this work incorporates client identifier information into the gating decision; Second, to reduce the difficulty of training MoE expert models on non-IID data, which could lead to model degradation, a variational lower bound is leveraged to derive a unified global variational approximation, which replaces the true posterior distribution of all clients during training. [131] proposes an improved model fine-tuning method, where a general model is first trained and each client only fine-tunes the shallow parameters in the general model. Additionally, for complex high-dimensional data, this work uses partial model parameters to extract features from the raw data before feeding them into the gating function to make inputs distinguishable. [132] first trains local models on each client and then uses the clustering method IFCA [133] to iteratively cluster all clients, generating a global cluster model for each cluster. Finally, a gating mechanism is used to combine the outputs of the global cluster models and the local models to produce the final result. [134] also introduces a MoE layer on the server side but dynamically determines the number of top-K experts to retain based on the client’s resource conditions during routing, fully utilizing limited computational resources.
In addition to the discussion on heterogeneity, another significant challenge in federated learning is communication overhead. Each round of communication between decentralized servers incurs substantial costs. Generally, there are two approaches to address this issue. The first approach is to reduce the amount of information that needs to be communicated in each round by transmitting only critical model parameters. For example, the number of experts in the locally trained MoE layer can be reduced, discarding some expert parameters that contribute less to the output in exchange for improved communication efficiency [135, 134]. Alternatively, only the most important MoE models can be selected for transmission. The second approach is to reduce the number of communication rounds, such as in One-shot Federated Learning (OFL), where typically only a single round of communication is performed during training to aggregate client models on the central server. After this, clients no longer participate in subsequent optimization updates. However, many popular OFL methods [136, 137] employ lossy aggregation, where aggregating all local models into a single global model results in knowledge loss, especially when model heterogeneity is strong [266]. In contrast, [138] proposes to treat all local models as expert models and train a specialized lightweight gating function using additional data to aggregate these experts. This approach effectively preserves the knowledge of local models while reducing communication overhead.
@@ IV Theory
Although the MoE model has a long history and achieved great empirical success in practice during the past few years [267, 16], the theory behind it is largely under-explored, especially in deep neural networks. Recently, some studies have initiated the attempt to build theoretical understandings of MoE. In this section, we will delve into these studies, aiming to provide the readers a basic idea of the theory development in this area.
@@
With the increasingly broad applications of MoE in many problems, the theoretical studies of MoE start to emerge beyond the general machine learning setup. [152] explores the usage of MoE models in handling multilevel data while building the theoretical understanding therein. Notably, multilevel data is very common in practical applications [268, 269], such as hierarchical data with a nested structure where different levels are correlated. Ignoring the dependencies within the data can lead to spurious, misleading, or biased clustering and prediction outcomes [270]. To better capture the dependencies, [152] proposes a Mixed MoE to handle multi-level data, which we define as MeMoE in this work. This approach not only introduces random effects into the gating function, making the computed gating values depend not only on input variables but also on random effects. This helps the model better capture random effects in multi-level data when dynamically adjusting expert weights. At the same time, to simplify the model, the expert functions are separated from the random effects, and the probability distribution of the expert model no longer depends on input variables but only on output variables. Under certain regularity conditions, this paper proves that the MeMoE is dense in the sense of weak convergence for any continuous mixed-effects model. In the special case of hierarchical multi-level data, this paper further proves that MeMoE can approximate the complex dependency structures of random effects between different factor levels in multi-level data.
@@ V Applications
In this section, we investigate the applications of MoE in two broad application domains, computer vision (CV) and natural language processing (NLP). By examining the influence of MoE in tasks ranging from image classification to machine translation, we highlight how this architectures enhance specialization, scalability, and adaptability, offering insights into their practical impact and future potential across diverse real-world scenarios.